# Word problems and embedding-obstructions in cellular automata groups on groups

### v2

@@ Abstract
We study groups of reversible cellular automata, or CA groups, on groups. More generally, we consider automorphism groups of subshifts of finite type on groups. It is known that word problems of CA groups on virtually nilpotent groups are in co-NP, and can be co-NP-hard. We show that under the Gap Conjecture of Grigorchuk, their word problems are PSPACE-hard on all other groups. On free and surface groups, we show that they are indeed always in PSPACE. On a group with co-NEXPTIME word problem, CA groups themselves have co-NEXPTIME word problem, and on the lamplighter group (which itself has polynomial-time word problem) we show they can be co-NEXPTIME-hard. We show also two nonembeddability results: the group of cellular automata on a non-cyclic free group does not embed in the group of cellular automata on the integers (this solves a question of Barbieri, Carrasco-Vargas and Rivera-Burgos); and the group of cellular automata in dimension $D$ does not embed in a group of cellular automata in dimension $d$ if $D\geq 3d+2$ (this solves a question of Hochman).
@@ 1.1 Complexity-theoretic results
Appendix A gives the (mostly standard) complexity-theoretic definitions used in the statements.
Our first theorem is barely new, and is stated mainly for context. The hard was shown essentially in [29] (it proves the $\mathbb{Z}$-case) and only a short additional discussion is needed here (see Corollary 8.1). The co-NP algorithm is essentially from [8, 23].
@@
Let $G$ be an f.g. group with have polynomial growth, and let $X\subset A^{G}$ be any subshift of finite type with language in NP. Then every finitely-generated subgroup of $\mathrm{Aut}(X)$ has word problem in co-NP. Furthermore, for any infinite f.g. group $G$ with polynomial growth, $\mathrm{Aut}(A^{G})$ has a f.g. subgroup with NP-complete word problem.
@@
We immediately note that the larger $\beta$ is, the more this conjecture is claiming, and that for large $\beta$ it states too much. Indeed, the conjecture is not true for e.g. $\beta>\log_{32}31=\gamma$, as the Grigorchuk group has growth $O(e^{n^{\gamma}})$ and is not virtually nilpotent [16] (much more is known about the growth of the Grigorchuk group [14]). For our purposes, it would suffice that this conjecture is true for any $\beta>0$.
@@
In several papers [31, 29, 9, 32], the author (with various coauthors) has studied a specific subgroup of $\mathrm{Aut}(A^{G})$ called $\mathrm{PAut}_{G}(A)$, which is generated by partial shifts (which split the aphabet into a Cartesian product and then shift one of the resulting tracks only), and symbol permutations (permuting the symbols in cells individually). This group is somewhat mysterious, but it provides a reasonably canonical f.g. subgroup, where many behaviors of the typically non-f.g. group $\mathrm{Aut}(A^{G})$ can already be seen.
@@
The main new idea needed is the use of fast information propagation along trees. Specifically, we can also work with events of exponential (or otherwise fast-growing) support, by using the parallelism of cellular automata. Unlike in Barrington’s theorem, it seems we cannot do this while preserving a perfect separation of control and permuted data. The technique we use for this, and in fact in all our constructions, is something we call dyadically accelerated funneling.
@@
For example $X=A^{G}$ satisfies the assumptions on $X$. We give more generally a sufficient geometric condition for such an upper bound in terms of “polynomial splitting schemes” for balls. We leave it as an open question what groups admit such schemes. One interesting question is whether they exist for small cancellation groups or hyperbolic groups with one-dimensional boundary.
Two remarks are in order. First, note that the word problem of $\mathrm{PAut}_{G}(A)$ contains the word problem of $G$, so of course if $G$ itself has PSPACE-hard word problem (or harder) then the theorem is trivial. The word problem is decidable in polynomial time in a large class of groups, and for example for a free group or a surface group it can be solved by Dehn’s algorithm [12, 7].
Second, it is known that automata groups (whose definition we omit) can have PSPACE-complete word problems [36] (and they always have PSPACE word problem [35]). Barrington’s theorem is a key component of the proof in [36], and it would be interesting to know how related the proofs are.
@@
On the lamplighter group $\mathbb{Z}_{2}\wr\mathbb{Z}$, for every full shift on at least two letters, the automorphism group has a finitely-generated subgroup with co-NEXPTIME-complete word problem. For any alphabet $A$ with sufficiently many prime factors, we can use the group $\mathrm{PAut}(A)$.
@@
We mention by analog that the domino problem, whether a given SFT is nonempty, is known to be undecidable on $\mathbb{Z}_{2}\wr\mathbb{Z}$, $\mathrm{BS}(1,n)$ [4].
Conjecture 1.2.
Word problems of CA groups can be co-NEXPTIME-complete on a large class of groups.
@@
It seems that in most cases, when the word problem was solved for a group of general interest, it ended up being solvable in polynomial time. In fact, we only know the following cases where the word problem ended up being complete for a class conjecturally not solvable in polynomial time:
@@
In the second item, the group explicitly used is $\mathrm{PAut}(A)$. One may wonder if this group is really natural: the generators are natural and study of the group emerged from questions unrelated to computability or complexity, but we do not even know, for example, whether it is decidable whether a given $f\in\mathrm{Aut}(A^{\mathbb{Z}})$ belongs to this group. However (unlike in e.g. Theorem C of the present paper), the proof in [29] in fact only uses the subgroup of $\mathrm{PAut}(A,B)$ where the $B$-track cannot affect the $A$-track in symbol permutations, which is a tamer and more natural group in many ways (see e.g. [31]).
A final remark we make here is that in each of the three items above, it is only conjectural (even if widely-believed) that there is no polynomial-time algorithm for the word problem, as PSPACE $\neq$ P is not a proven fact. However, Theorem E unconditionally shows that there is no polynomial-time algorithm for the word problem of $\mathrm{PAut}(A)$ on the lamplighter group, and for alphabets $A$ with sufficiently many prime factors.
@@ 1.2 Non-embeddability results
These are based on an observation that we believe is entirely new (although in retrospect, rather obvious), namely that residual finiteness of $\mathrm{Aut}(A^{\mathbb{Z}^{d}})$ admits a quantitative variant when working with finitely-generated subgroups, based on the fact that nontrivial automorphism with radius $n$ acts nontrivially on $n$-periodic points. See Lemma 4.1. This idea bears resemblance to the use of simple groups in [34, 13, 32], although we do not know if there is a direct connection.
@@
Suppose that $D>3d+1$. Then $\mathrm{Aut}(A^{\mathbb{Z}^{D}})$ has a finitely-generated subgroup that does not embed in $\mathrm{Aut}(B^{\mathbb{Z}^{d}})$ for any nontrivial alphabets $A,B$.
So, for example, $\mathrm{Aut}(\{0,1\}^{\mathbb{Z}^{5}})$ does not embed in $\mathrm{Aut}(\{0,1\}^{\mathbb{Z}})$, and $\mathrm{Aut}(\{0,1\}^{\mathbb{Z}^{8}})$ does not embed in $\mathrm{Aut}(\{0,1\}^{\mathbb{Z}^{2}})$. There are no previously known nonembeddability results for the groups $\mathrm{Aut}(A^{\mathbb{Z}^{d}})$, indeed previously the only embeddability obstructions that have been observed or applied are (qualitative) residual finiteness and complexity of the word problem.
@@
It should be possible to optimize the proof to get nonembeddings for $D\geq 3d+1$, as our proof “wastes” one dimension, in a sense. It is an interesting question how far this can be pushed, in particular whether all cases $D\geq d+1$ can be covered.
As noted in [2], the groups $\mathrm{Aut}(A^{\mathbb{Z}^{d}})$ and $\mathrm{Aut}(A^{\mathbb{Z}^{D}})$ are non-isomorphic for $d\neq D$, since their centers are the respective shift groups $\mathbb{Z}^{D}$ and $\mathbb{Z}^{d}$. This is a result of Hochman [19] (generalizing Ryan’s theorem [27], and further generalized by Barbieri, Carrasco-Vargas and Rivera-Burgos in [2]). However, it seems impossible to obtain any results about embeddability using this kind of method.
@@
Suppose that $D>3d+1$, $\mathcal{G}_{d}=\mathrm{Aut}(A^{\mathbb{Z}^{d}}),\mathcal{G}_{D}=\mathrm{Aut}(B^{\mathbb{Z}^{D}})$ for any nontrivial alphabets $A,B$. Then $\mathcal{G}_{D}$ has a finitely-generated subgroup such that $\mathcal{G}_{D}/K$ does not embed in $\mathcal{G}_{d}/K^{\prime}$ for any central subgroups $K\leq\mathcal{G}_{D},K^{\prime}\leq\mathcal{G}_{d}$.
@@
We solve this question completely.
@@
Let $G$ be a f.g. group containing a free submonoid on two generators, and let $A$ be any nontrivial alphabet. Then $\mathrm{Aut}(A^{G})$ does not embed into $\mathrm{Aut}(B^{\mathbb{Z}^{d}})$ for any alphabet $B$ and $d\in\mathbb{N}$.
@@ Acknowledgements
We thank Ilkka Törmä for working through the proof of Theorem 4.1 in full detail. Of course, any mistakes are mine. The connection between the Gap Conjecture and a PSPACE frontier originated from the study of another computational model with Ilkka Törmä (apart from the construction of markers). We thank Sebastián Barbieri for sharing Question 1.2, and Scott Schmieding for several inspiring discussions. We thank Laurent Bartholdi for helping streamline the surface group argument (Lemma 8.2), and other useful comments. We thank Michael Hochman for asking about Theorem G.
@@ 2 Definitions, conventions and some standard observations
In this section, we explain our definitions and conventions. This does not include complexity-theoretic definitions, which are quite lengthy. We have instead written these in an appendix, which also contains some basic results on complexity theory.
The appendix is aimed mainly at readers with zero knowledge and infinite intelligence, and some readers may find it helpful to also consult a textbook [1]. Readers familiar with complexity theory will not need to read Section A.1 and Section A.2 in the appendix, but may find it helpful to check out Lemma A.3 (an explicit tree formulation of a computation of an alternating Turing machine in terms of two nondeterministic Turing machines). We also give a proof of a version of Barrington’s theorem in Section A.3. A reader familiar with previous work of the author may mostly skip Section A.3, while for others (even readers familiar with Barrington’s theorem) we recommend taking a look, as we use similar notation and arguments in automorphism groups.
2.1 Assorted definitions
@@
For words $u,v$ we denote the free monoid operation, i.e. concatenation, by $uv$. The empty word (unique word of length 00) is written $\epsilon$, and we write $A^{+}=A^{*}\setminus\{\epsilon\}$. A set of words $W\subset A^{*}$ is mutually unbordered if, whenever $u,v\in W$, $i,j\in\mathbb{Z}$, and $x\in A^{\mathbb{Z}}$ satisfies $x|[i,i+|u|-1]=u$, $x|[j,j+|v|-1]=v$, we have $[i,i+|u|-1]\cap[j,j+|v|-1]=\emptyset$, unless $i=j\wedge u=v$. A word $w$ is unbordered if $\{w\}$ is mutually unbordered.
@@
In a metric space $X$, we say $Y\subset X$ is $r$-dense if every $r$-ball intersects $Y$. We say $Y$ is $r$-sepated if $d(y,y^{\prime})<r$ for $y,y^{\prime}\in Y$ implies $y=y^{\prime}$. The following is immediate in finite metric spaces (in general, axiom of choice is needed):

### v3

@@ Abstract
We study groups of reversible cellular automata, or CA groups, on groups. More generally, we consider automorphism groups of subshifts of finite type on groups. It is known that word problems of CA groups on virtually nilpotent groups are in co-NP, and can be co-NP-hard. We show that under the Gap Conjecture of Grigorchuk, their word problems are PSPACE-hard on all other groups. On free and surface groups, we show that they are indeed always in PSPACE. On a group with co-NEXPTIME word problem, CA groups themselves have co-NEXPTIME word problem, and on the lamplighter group (which itself has polynomial-time word problem) we show they can be co-NEXPTIME-hard. We show also nonembeddability results: the group of cellular automata on a non-cyclic free group does not embed in the group of cellular automata on the integers (this solves a question of Barbieri, Carrasco-Vargas and Rivera-Burgos); and the group of cellular automata in dimension $D$ does not embed in a group of cellular automata in dimension $d$ if $D>d$ (this solves a question of Hochman).
@@ 1.1 Complexity-theoretic results
Appendix B gives the (mostly standard) complexity-theoretic definitions used in the statements.
Our first theorem is barely new, and is stated mainly for context, as it was shown essentially in [28]. Namely [28] proves the $\mathbb{Z}$-case, and only a short additional discussion is needed here (see Corollary 9.2). The co-NP algorithm is essentially from [10, 24].
@@
Let $G$ be an f.g. group with polynomial growth, and let $X\subset A^{G}$ be any subshift of finite type with language in NP. Then every finitely-generated subgroup of $\mathrm{Aut}(X)$ has word problem in co-NP. Furthermore, for any infinite f.g. group $G$ with polynomial growth, $\mathrm{Aut}(A^{G})$ has a f.g. subgroup with NP-complete word problem.
@@
We immediately note that this is a parametrized family of conjectures, and the larger $\beta$ is, the more it is claiming. For large $\beta$, it claims too much: the conjecture is not true for e.g. $\beta>\log_{32}31=\gamma$, as the Grigorchuk group has growth $O(e^{n^{\gamma}})$ and is not virtually nilpotent [18] (much more is known about the growth of the Grigorchuk group [16]). For our purposes, it would suffice that this conjecture is true for a single (possibly very small) $\beta>0$.
@@
In several papers [30, 28, 11, 31], the author (with various coauthors) has studied a specific subgroup of $\mathrm{Aut}(A^{G})$ called $\mathrm{PAut}_{G}(A)$, which is generated by partial shifts (which split the aphabet into a Cartesian product and then shift the configuration on one of the resulting components), and symbol permutations (permuting the symbols in cells individually). This group is somewhat mysterious, but it provides a reasonably canonical f.g. subgroup, where many behaviors of the typically non-f.g. group $\mathrm{Aut}(A^{G})$ can already be seen.
@@
The main new idea needed is the use of fast information propagation along trees. Specifically, we can also work with events of exponential (or otherwise fast-growing) support, by using the parallelism of cellular automata. Unlike in Barrington’s theorem, it seems we cannot do this while preserving a perfect separation of control and permuted data. The technique we use for this, and in fact in all our constructions, is something we call ripple catching.
@@
For example $X=A^{G}$ satisfies the assumptions on $X$. We give more generally a sufficient geometric condition for such an upper bound in terms of “polynomial splitting schemes” for balls.
Two remarks are in order. First, note that the word problem of $\mathrm{PAut}_{G}(A)$ contains the word problem of $G$, so of course if $G$ itself has PSPACE-hard word problem (or harder) then the PSPACE-hardness theorem is trivial. The word problem is, however, decidable in polynomial time in a large class of groups, and for example for a free group or a surface group it can be solved by Dehn’s algorithm [14, 7].
Second, it is known that automata groups (whose definition we omit) can have PSPACE-complete word problems [35] (and they always have PSPACE word problem [34]). Barrington’s theorem is also a key component of the proof in [35].
@@
On the lamplighter group $\mathbb{Z}_{2}\wr\mathbb{Z}$, for every full shift on at least two letters, the automorphism group has a finitely-generated subgroup with co-NEXPTIME-complete word problem. For some alphabets $A$, we can use the group $\mathrm{PAut}(A)$.
@@
It seems that in most cases, when the complexity of the word problem of a group of general interest is known, it is known to be solvable in polynomial time. In fact, we only know the following cases where the word problem ended up being complete for a class conjecturally not solvable in polynomial time:
@@
In the second item, the group explicitly used is $\mathrm{PAut}(A)$. One may wonder if this group is really natural: the generators are natural and study of the group emerged from questions unrelated to computability or complexity, but we do not even know, for example, whether it is decidable whether a given $f\in\mathrm{Aut}(A^{\mathbb{Z}})$ belongs to this group. However (unlike in e.g. Theorem C of the present paper), the proof in [28] in fact only uses the subgroup of $R(A,B)$ of $\mathrm{PAut}(A\times B)$ where the $B$-track cannot affect the $A$-track in symbol permutations, which is a tamer and more natural group in many ways, and for example is elementarily amenable (see [30]).
A final remark we make here is that in each of the three items above, it is only conjectural (even if widely-believed) that there is no polynomial-time algorithm for the word problem, as PSPACE $\neq$ P is not a proven fact. However, Theorem E unconditionally shows that there is no polynomial-time algorithm for the word problem of $\mathrm{PAut}(A)$ on the lamplighter group, and for some alphabets $A$.
@@ 1.2 Non-embeddability results
These are based on an observation that we believe is entirely new (although in retrospect, rather obvious), namely that residual finiteness of $\mathrm{Aut}(A^{\mathbb{Z}^{d}})$ admits a quantitative variant when working with finitely-generated subgroups, based on the fact that a nontrivial automorphism with radius $n$ acts nontrivially on points with period $O(n)$ in each cardinal direction. See Lemma 4.1. This idea bears resemblance to the use of simple groups in [33, 15, 31].
@@
Suppose that $D>d$. Then $\mathrm{Aut}(A^{\mathbb{Z}^{D}})$ has a finitely-generated subgroup that does not embed in $\mathrm{Aut}(B^{\mathbb{Z}^{d}})$ for any nontrivial alphabets $A,B$.
There are no previously known nonembeddability results for the groups $\mathrm{Aut}(A^{\mathbb{Z}^{d}})$, indeed previously the only embeddability obstructions that have been observed or applied even in the case of $\mathrm{Aut}(A^{\mathbb{Z}})$ are residual finiteness (i.e. all subgroups must be residually finite) and complexity of the word problem (i.e. all f.g. subgroups must have coNP word problem).
@@
As noted in [2], the groups $\mathrm{Aut}(A^{\mathbb{Z}^{d}})$ and $\mathrm{Aut}(B^{\mathbb{Z}^{D}})$ are non-isomorphic for $d\neq D$, since their centers are the respective shift groups $\mathbb{Z}^{D}$ and $\mathbb{Z}^{d}$. This is a result of Hochman [21] (generalizing Ryan’s theorem [27], and further generalized by Barbieri, Carrasco-Vargas and Rivera-Burgos in [2]). However, it seems impossible to obtain any results about embeddability using this kind of method.
@@
Suppose that $D>d$, $\mathcal{G}_{d}=\mathrm{Aut}(A^{\mathbb{Z}^{d}}),\mathcal{G}_{D}=\mathrm{Aut}(B^{\mathbb{Z}^{D}})$ for any nontrivial alphabets $A,B$. Then $\mathcal{G}_{D}$ has a finitely-generated subgroup such that $\mathcal{G}_{D}/K$ does not embed in $\mathcal{G}_{d}/K^{\prime}$ for any central subgroups $K\leq\mathcal{G}_{D},K^{\prime}\leq\mathcal{G}_{d}$.
@@
We solve this question as well.
@@
Let $G$ be a f.g. group with at least stretched exponential growth. Then $\mathrm{Aut}(A^{G})$ has a finitely generated subgroup that does not embed into $\mathrm{Aut}(B^{\mathbb{Z}^{d}})$ for any alphabet $B$ and $d\in\mathbb{N}$.
@@ Acknowledgements
We thank Ilkka Törmä for working through our first proof of Theorem 4.6 in full detail. Of course, any mistakes are mine. The connection between the Gap Conjecture and a PSPACE frontier originated from the study of another computational model with Ilkka Törmä (apart from the construction of markers). We thank Sebastián Barbieri for sharing Question 1.3, and Scott Schmieding for several inspiring discussions. We thank Laurent Bartholdi for helping to streamline the surface group argument (Lemma 9.7), and other useful comments. We thank Michael Hochman for asking about Theorem G.
@@ 2 Definitions, conventions and some standard observations
In this section, we explain our definitions and conventions. In the appendix, we have included an index (Section A), which may help keep track of our notations.
Complexity-theoretic definitions are not found in this section, as they are quite lengthy, and mostly standard. We have instead written these in an appendix, which also contains some basic results on complexity theory. The reader may find it helpful to also consult a textbook such as [1].
Readers familiar with complexity theory will not need to read Section B.1 and Section B.2 in the appendix, but may find it helpful to check out Lemma B.4 (an explicit tree formulation of a computation of an alternating Turing machine in terms of two nondeterministic Turing machines). We also give a proof of a version of Barrington’s theorem in Section B.3. A reader familiar with previous work of the author may mostly skip Section B.3, while for others (even readers familiar with Barrington’s theorem) we recommend taking a look, as we use similar notation and arguments in automorphism groups.
2.1 Preliminaries
@@
By $\bar{X}$ we mean the complement of the set $X$ when the containing set is clear from context. By $A\sqcup B$ we denote a union of two sets $A,B$, when $A\cap B=\emptyset$ (i.e. this is a domain-restriction of the union operation). In a cartesian product $A\times B$, we refer to $A$ and $B$ as components, but especially in the context of sets $(A\times B)^{G}$, we refer to $A$ and $B$, as well as the projections $A^{G}$ and $B^{G}$, as tracks. Intervals $[i,j]$ are interpreted on $\mathbb{Z}$, i.e. $[i,j]=\{k\in\mathbb{Z}\;|\;i\leq k\leq j\}$, unless explicitly specified otherwise. By $\mathbb{Z}_{k}$ we denote the cyclic group with $k$ elements. By $1_{G}$ we denote the identity element of a group $G$.
In many cases, we will have either a graph $G$ or a configuration $x$, and on its nodes $u\in G$ or positions $x_{g}$, we have colors (or symbols). We often think of the graph or configuration as being fixed, and speak of the color at $u$ or $g$. The color in turn is often a Cartesian product of the form $C\times B$. Then $B$-value of a node $u$ or element $g$ refers to the $B$-component of $x_{g}$ or of the color of $u$, and similarly for $C$.
For a function $g:\mathbb{N}\to\mathbb{R}$ write $O(g)$ (resp. $\Omega(g)$) for functions $f$ such that for some $C>0$, $f(n)\leq Cg(n)$ (resp. $f(n)\geq Cg(n)$) for large enough $n$. Write $\Theta(g)$ for the intersection. It is customary to use the notation $f=O(g),f=\Omega(g),f=\Theta(g)$ instead of inclusion.
For nondecreasing functions $f,g:R\to\mathbb{R}$ with $R\subset\mathbb{R}$, write $f\lesssim g$ if there exist $a,b>0$ such that for some $r_{0}$, for all $r_{0}\leq r\in R$ we have $br\in R$ and $f(r)\leq ag(br)$. Write $f\gtrsim g\iff g\lesssim f$. If $f\lesssim g$ and $f\gtrsim g$, we write $f\asymp g$ and say that $f,g$ are asymptotically equal. Note that all functions in $f(\Theta(n))$ are asymptotically equal. The term growth refers to a class of functions closed under asymptotic equality.
Lemma 2.1.
If $f(n\log^{d}n)\lesssim g(n)$, then $f(n)\lesssim g(n/\log^{d}n)$. Similarly for $\gtrsim$.
Proof.
Setting $m=n/\log^{d}n$ we have
$m\log^{d}m=(n/\log^{d}n)\log^{d}(n/\log^{d}n)=n-n\log^{d}\log^{d}n/\log^{d}n=\Theta(n)$
Thus if $f(n\log^{d}n)\lesssim g(n)$, we have
$f(n)\asymp f(\Theta(m\log^{d}m))\lesssim g(m)=g(n/\log^{d}n).$
The case of $\gtrsim$ is similar. ∎
@@
For words $u,v$ we denote the free monoid operation, i.e. concatenation, by $uv$. The empty word (unique word of length 00) is written $\epsilon$, and we write $A^{+}=A^{*}\setminus\{\epsilon\}$. A set of words $W\subset A^{*}$ is mutually unbordered if, whenever $u,v\in W$, $i,j\in\mathbb{Z}$, and $x\in A^{\mathbb{Z}}$ satisfies $x|[i,i+|u|-1]=u$, $x|[j,j+|v|-1]=v$, we have $[i,i+|u|-1]\cap[j,j+|v|-1]=\emptyset$, unless $i=j$ and $u=v$. A word $w$ is unbordered if $\{w\}$ is mutually unbordered.
@@
In a metric space $X$, we say $Y\subset X$ is $r$-dense if every $r$-ball intersects $Y$. We say $Y$ is $r$-separated if $d(y,y^{\prime})<r$ for $y,y^{\prime}\in Y$ implies $y=y^{\prime}$. The following is immediate in finite metric spaces (in general, axiom of choice is needed):